One month Lisa rented
2
movies and
6
video games for a total of
$36
. The next month she rented
5
movies and
3
video games for a total of
$27
. Find the rental cost for each movie and each video game.

Found 2 solutions by FrankM, stanbon:
Answer by FrankM(1040)

2M+6V=36
5M+3V=27
10M+6V=54 (eq2 *2)
2M+6V=36
8M=18
M=18/8 or 2.25
2(2.25)+6V=36
4.50+6V=36
6V=31.50
V=5.25
Movies cost $2.25 Video games $5.25

Answer by stanbon(75887)

One month Lisa rented
2
movies and
6
video games for a total of
$36
----
2m + 6v = 36 dollars
----------------------------------
. The next month she rented
5
movies and
3
video games for a total of
$27
5m + 3v = 27 dollars
==============================
. Find the rental cost for each movie and each video game.
2m + 6v = 36 dollars
5m + 3v = 27 dollars
--------------
Modify for elimination::
10m + 30v = 5*36
10m + 6v = 2*27
---------------------
Subtract and solve for "v"::
24v = 126
----
each video game costs 126/14 = $9.00
-------------
Solve for "m":
2m + 6v = 36
2m + 54 = 36
Note:: This results in a negative cost for movies.
Please check your post.
=========
Cheers,
Stan H.
--------
=============